OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)? 4x - y = 1 4x + y = 1 x - 4y = -13 4x - y = -7

OpenStudy (anonymous):

any ideas? first you need the slope

OpenStudy (anonymous):

What's the formula for finding the slope?

OpenStudy (anonymous):

change in \(y\) over the change in \(x\)

OpenStudy (anonymous):

sometimes written \(\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

Oh yeah. Okay give me a second to to this.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got 4.

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

Now what is the next step?

OpenStudy (anonymous):

you have a choice you can use the point-slope formula to find the equation of the line, or since you only have 4 choices you can pick the one that has slope of 4 in fact we should probably use the first method, \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

How would I set that up?

OpenStudy (anonymous):

oh in fact we have to use the first method, since there are two lines with slope 4 make the number replacements in the formula \[y-y_1=m(x-x_1)\] with \(x_1=1,y_1=11, m=4\)

OpenStudy (anonymous):

you should get \[y-11=4(x-1)\]

OpenStudy (anonymous):

>.< Ah, I hate this part.

OpenStudy (anonymous):

yeah i know, but the steps are always the same

OpenStudy (anonymous):

just like yesterday

OpenStudy (anonymous):

So minus the 1 from 11?

OpenStudy (anonymous):

multiply out using the distributive law to get \[y-11=4x-4\] no distribute first that is always the first step

OpenStudy (anonymous):

Oh yeah. Okay so. y-11=4x-4

OpenStudy (anonymous):

\[(y-11)/(x-1)=(11-3)/(1+1)\] \[(y-11)/(x-1)=4\] \[4x-y=-7\]

OpenStudy (anonymous):

make sure to use the distributive law to get rid of the parentheses

OpenStudy (anonymous):

then you can add 11 to both sides \[y=4x+7\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

we had \[y-11=4x-4\] and now we want to write it in standard form

OpenStudy (anonymous):

first we can add 11 to both sides to get \[y=4x+7\] is that ok?

OpenStudy (anonymous):

So then we would get 4x - y = -7?

OpenStudy (anonymous):

this is the "slope intercept" form from last night

OpenStudy (anonymous):

then we can put all the variables on one side, the number on the other

OpenStudy (anonymous):

Oh I see. I'm very slow at learning. Takes me a million times to do something before it's pinned in my small brain for good.

OpenStudy (anonymous):

subtract \(y\) from both sides and subtract 7 from both sides and get the answer you wrote

OpenStudy (anonymous):

of course it takes practice. you were not born knowing this

OpenStudy (anonymous):

but the steps are always the same so if you do several it will become easier

OpenStudy (anonymous):

we can try another if you like

OpenStudy (anonymous):

Thanks for your help (: We can practice some tomorrow. I'm doing a practice assignment that is due soon.

OpenStudy (anonymous):

good luck